Please help me solve this word problem:
Gustav invested all of his $3,000 savings: a part in annual rate of 6% and the rest at an annual rate of 9%. His annual income from both investments is equivalent to 8% of his total investment. How much is invested at 9%?

Answer by rfer(16322)

x= 6% investment
3000-x=9% investment
.06x+.09(3000-x)=.08(3000)
.06x-.09x=240-270
-.03x=-30
x=$1000
3000-x=$2000